OpenStudy (anonymous):

Integrate. sinx+secx/tanx I think I messed up some where because I get sinx-cscx.

OpenStudy (lgbasallote):

\[\int \frac{\sin x + \sec x}{\tan x} dx?\]

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

\[\int \frac{\sin x }{\tan x} dx + \frac{\sec x}{\tan x}dx\] this is what you did right?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

there's an integral for sec x/tan x of course

OpenStudy (lgbasallote):

\[\large \int \frac{\sin x}{\frac{\sin x}{\cos x}}dx + \int \frac{\frac{1}{\cos x}}{\frac{\sin x}{\cos x}}dx\] is this what you did next?

OpenStudy (anonymous):

Yes

OpenStudy (lgbasallote):

then what did you do?

OpenStudy (anonymous):

Oh, I put in a sin^2 at the bottom, that's why I messed up.

OpenStudy (lgbasallote):

so what did you get in this one?

OpenStudy (anonymous):

I got sinx+ln|cscx-cotx|+c

OpenStudy (lgbasallote):

hmm

OpenStudy (lgbasallote):

seems right congrats